# Comparing the information content of probabilistic representation spaces

### v2

@@ Abstract
Probabilistic representation spaces convey information about a dataset, and to understand the effects of factors such as training loss and network architecture, we seek to compare the information content of such spaces. However, most existing methods to compare representation spaces assume representations are points, and neglect the distributional nature of probabilistic representations. Here, instead of building upon point-based measures of comparison, we build upon classic methods from literature on hard clustering. We generalize two information-theoretic methods of comparing hard clustering assignments to be applicable to general probabilistic representation spaces. We then propose a practical method of estimation that is based on fingerprinting a representation space with a sample of the dataset and is applicable when the communicated information is only a handful of bits. With unsupervised disentanglement as a motivating problem, we find information fragments that are repeatedly contained in individual latent dimensions in VAE and InfoGAN ensembles. Then, by comparing the full latent spaces of models, we find highly consistent information content across datasets, methods, and hyperparameters, even though there is often a point during training with substantial variety across repeat runs. Finally, we leverage the differentiability of the proposed method and perform model fusion by synthesizing the information content of multiple weak learners, each incapable of representing the global structure of a dataset. Across the case studies, the direct comparison of information content provides a natural basis for understanding the processing of information.
@@ 1 Introduction
The comparison of representation spaces is a problem that has received much attention, particularly as a route to a deeper understanding of information processing systems (Klabunde et al., 2023; Mao et al., 2024; Huh et al., 2024). Existing methods are applicable to point-based representation spaces, including centered kernel alignment (CKA) (Kornblith et al., 2019) and representational similarity analysis (RSA) (Kriegeskorte et al., 2008). For representation spaces whose citizens are probability distributions, such as in variational autoencoders (VAEs) or in biological systems with inherent stochasticity, failure to account for the distributed nature of representations can miss important aspects of the relational structure between data points (Duong et al., 2023). There are surprisingly few options for comparing spaces that fully account for the distributional nature of representations (Klabunde et al., 2023).
One of the only existing methods, stochastic shape metrics (Duong et al., 2023), extends a point-based approach that quantifies the difficulty of aligning one representation space to another through a specified set of allowed transformations (Williams et al., 2021). Here we argue that an information theoretic perspective is most natural when considering the role of a representation space as an intermediate stage of processing in a deep neural network. To be specific, a probabilistic representation space communicates certain information about the data, and we propose to compare two such spaces via quantities related to the mutual information between them. A consequence of focusing on information content is that the spaces can differ in dimensionality, in contrast to shape metrics (Duong et al., 2023), and in whether they are discrete or continuous.
We take as a motivating example the task of unsupervised disentanglement, whose goal is to break information about a dataset into useful pieces. What qualifies as useful depends on the context: although statistical independence between the pieces of information is the predominant desideratum (Locatello et al., 2019), interpretability of the factors can be prioritized (Molnar, 2022) and will clash with independence when correlations are present (Träuble et al., 2021). Other useful properties proposed for disentanglement are compositionality under group operations (Higgins et al., 2018; Balabin et al., 2023) and performance enhancement on downstream tasks (Van Steenkiste et al., 2019).
A significant challenge in disentanglement research is evaluation, which relies on assessing the fragmentation of information about the dataset into channels of a trained model. Models are predominantly evaluated against ground truth factors of variation on overly synthetic datasets (Chen et al., 2018; Kim & Mnih, 2018; Eastwood & Williams, 2018; Hsu et al., 2023). It remains unclear how to evaluate disentanglement when datasets are not accompanied by known generative factors, hindering our understanding of how disentanglement methods work on more realistic datasets. Methods of unsupervised evaluation have been proposed based on a notion of model centrality (Duan et al., 2020; Lin et al., 2020): models that split information similarly to other models are presumed disentangled. Yet, relatedness has failed to account for the probabilistic nature of the representation spaces of the models. Instead, posterior distributions are reduced to points by taking the means of posterior distributions or sampling, and then point-based comparisons such as correlation are applied (Duan et al., 2020).
@@
In this work, we treat probabilistic representation spaces as soft clustering assignments of data, whereby partial distinguishability between data points is expressed by the overlap between posterior distributions. We generalize classic information-theoretic measures of similarity between hard clusterings (Fig. 1a) to be applicable to probabilistic representation spaces (Fig. 1b). By comparing the information content of learned representation spaces produced by multiple models in an ensemble, we study the effects of method, training progress, and dataset on information processing.
Our primary contributions are the following:
  1. 1.
    We generalize two classic measures that compare the information content of clusterings for the comparison of probabilistic representation spaces.
  2. 2.
    We propose a lightweight method to assess information content based on fingerprinting each representation space with the distinguishability of a sample from the dataset.
  3. 3.
    We demonstrate model fusion with the representation spaces from a set of weak VAEs, leveraging the differentiability of the method to maximize the similarity of a synthesis space with respect to the set.
@@
Figure 1: Similarity of generalized representation spaces. (a) A hard clustering assignment, such as the living/non-living distinction conveyed by clustering $V$, communicates certain information about the dataset (here, CIFAR-10 images). Comparing the information content of different clustering assignments enables comparative analyses between algorithms, model fusion, and benchmarking. (b) Here we extend measures for hard clusterings to be applicable to generalized representation spaces. There is no requirement for the dimensionality of the spaces to match, and hard clusterings can be compared to probabilistic spaces.
@@ 2 Related work
The proposed method builds upon classic means of comparing different cluster assignments (clusterings) of a dataset and generalizes them to compare the information content of probabilistic representation spaces. We then use the method to empirically study generative models designed to fragment information about a dataset in a learned latent space, i.e., for unsupervised disentanglement.
2.1 Similarity of clusterings and of representation spaces
@@
The capacity to compare transformations of data produced by different machine learning models enables ensemble learning, a deeper understanding of methodology, and benchmarking (Punera & Ghosh, 2007). Strehl & Ghosh (2002) used outputs of clustering algorithms to perform ensemble learning, based on a measure of similarity between clustering assignments that we will extend in this work: the normalized mutual information (NMI). Referred to as consensus clustering or ensemble clustering, efforts to combine multiple clustering outputs can leverage any of a variety of similarity measures (Wagner & Wagner, 2007; Vinh et al., 2009; 2010; Vega-Pons & Ruiz-Shulcloper, 2011; Huang et al., 2017). Another measure backed by information theory is the variation of information (VI), coined by Meilă (2003) for clusters but recognized as a metric distance between information sources at least twice before (Shannon, 1953; Crutchfield, 1990).
@@
Commonly referred to simply as ‘clustering’, hard clusterings assign every input datum to one and only one output cluster and have been generalized to multiple forms of soft clustering (Campagner et al., 2023). We focus specifically on fuzzy clustering, where membership is assigned by degree to multiple clusters and must sum to one for each datum (Zadeh, 1965; Dunn, 1973; Ruspini et al., 2019). Probabilistic representation spaces can be viewed as communicating a soft assignment over latent vectors for each datum, with degree of membership expressed by the posterior distributions. Extensions of hard clustering comparisons to fuzzy clusters have been proposed for measures that only indirectly assess information content by counting agreement of assignments, such as the Rand index (Punera & Ghosh, 2007; Hullermeier et al., 2011; D’Ambrosio et al., 2021; Andrews et al., 2022; Wang et al., 2022). Information-based measures have been extended for specific types of soft clusters (Campagner & Ciucci, 2019; Campagner et al., 2023) but none, to our awareness, for comparing the information content of fuzzy clusterings over a continuous representation space.
@@
A rich area of research compares point-based representation spaces via the pairwise geometric similarity of a common set of data points in the space (Kornblith et al., 2019; Hermann & Lampinen, 2020; Klabunde et al., 2023), building upon representational similarity analysis from neuroscience (Kriegeskorte et al., 2008). Stochastic shape metrics (Duong et al., 2023) are a probabilistic extension to a distance metric between two point-based representation spaces based on aligning one with another through prescribed transformations (Williams et al., 2021). Instead of using point-based methods as our starting point, we build upon hard clustering literature because of the extensive connections with information theory.
2.2 Unsupervised disentanglement
@@
Disentanglement is the problem of splitting information into useful pieces, possibly for interpretability, compositionality, or stronger representations for downstream tasks. Shown to be impossible in the fully unsupervised case (Locatello et al., 2019; Khemakhem et al., 2020), research has moved to investigate how to utilize weak supervision (Khemakhem et al., 2020; Sanchez et al., 2020; Vowels et al., 2020; Murphy et al., 2022) and incorporate inductive biases (Balabin et al., 2023; Chen et al., 2018; Rolinek et al., 2019; Zietlow et al., 2021; Hsu et al., 2024).
@@
A significant challenge for disentanglement is evaluation and model selection when no ground truth is available. When a dataset’s generative factors are available for evaluation, models are compared against an idealized disentanglement where the factors are whole and separate in latent dimensions (Chen et al., 2018; Kim & Mnih, 2018) and simply encoded (Eastwood & Williams, 2018; Hsu et al., 2023). However, methods of unsupervised model evaluation are few and limited. Duan et al. (2020) proposed that disentangled models are more similar to each other than are entangled ones, in an ensemble of repeats with different initializations. They evaluated the similarity between two models with an ad hoc function of dimension-wise similarity, which was computed as the rank correlation or the weights of a linear model between embeddings in different one-dimensional spaces. ModelCentrality (Lin et al., 2020) is similar: the model most central in an ensemble is taken to be the most disentangled, with similarity quantified with the FactorVAE score (Kim & Mnih, 2018), where one model’s embeddings serve as the labels for another model. PIPE (Estermann & Wattenhofer, 2023) is an unsupervised disentanglement metric that relies on a characterization of the posterior distributions in a single model rather than an ensemble. In this work we shift focus from assessing model similarity to channel similarity, under the premise that the fragmentation of information at the core of disentanglement is more naturally studied via the information fragments themselves.
@@ 3 Method
Our goal is to compare the information transmitted about a dataset by different representation spaces, and we will use the comparison of hard clustering assignments (e.g., the output of $k$-means) as a point of reference. Analogously to hard clustering (Fig. 1a), probabilistic representation spaces communicate a soft assignment over embeddings that is expressed by a probability distribution in the representation space for each data point (Fig. 1b).
While the proposed method can be applied to any representation space with probabilistic embeddings, we will focus primarily on variational autoencoders (VAEs) (Kingma & Welling, 2014). Let the random variable $X$ represent a sample $x\sim p(x)$ from the dataset under study. $X$ is transformed by a stochastic encoder parameterized by a neural network to a variable $U=f(X,\epsilon)$, where $\epsilon$ is a source of stochasticity. We will make use of two fundamental quantities in information theory (Cover & Thomas, 1999), the entropy of a random variable $H(Z)=\mathbb{E}_{z\sim p(z)}[-\log p(z)]$ and the mutual information between two random variables $I(Y;Z)=H(Y)+H(Z)-H(Y,Z)$. The encoder maps each datapoint $x$ to a posterior distribution in the latent space, $p(u|x)$, and the information passed into $U$ about $X$ is regularized by a variational upper bound (Alemi et al., 2017), the expected Kullback-Leibler (KL) divergence (Cover & Thomas, 1999) between the posterior distributions and an arbitrary prior $r(u)$: $\mathbb{E}_{x\sim p(x)}[D_{\text{KL}}(p(u|x)||r(u))]$. Commonly, the posteriors and the prior are parameterized as normal distributions with diagonal covariance matrices, which will facilitate many of the involved measurements but is not required for the method.
@@ 3.1 Comparing representation spaces as soft clusterings
Consider a hard clustering of data as communicating certain information about the data (Fig. 1a). By observing the cluster assignment $U$ instead of a sample $X$ from the dataset, information $I(X;U)$ will have been conveyed. For a hard clustering, every data point is assigned unambiguously to a cluster—i.e., $H(U|X)=0$—which makes the communicated information equal to the entropy of the clustering, $I(X;U)=H(U)$.
@@
Soft clustering generalizes hard clustering to allow each datum to have partial membership in multiple clusters, effectively communicating partial distinguishability between data points (Zadeh, 1965; Ruspini et al., 2019) (Fig. 1b). Here we view a probabilistic representation space as a soft clustering, where a datum’s posterior distribution $p(u|x)$ expresses a soft assignment over the points $u$.
The utility of NMI or VI instead of $I(U;V)$ to compare clustering assignments largely resides in the standardization of their values. For two identical hard clustering assignments, $U$ and $U^{\prime}$, $\text{NMI}(U,U^{\prime})=1$ and $\text{VI}(U,U^{\prime})=0$. For identical soft assignments—e.g., two copies of a VAE’s encoder—NMI will generally not be 1 and VI will not be zero. In contrast to hard clustering assignments, soft assignments include additional entropy resulting from uncertainty that is separate from the information communicated about the dataset, $H(U)=I(X;U)+H(U|X)$. The entropy terms in Eqns. 1 and 2 that ground the mutual information $I(U;V)$ thus include entropy that is irrelevant to the communicated information. We posit that a natural generalization of NMI and VI replaces the entropy of a clustering assignment with the mutual information between two copies of that assignment: $H(U)$ becomes $I(U;U^{\prime})$, and similarly for $V$.
The generalized forms become
$\text{NMI}(U,V)=\frac{I(U;V)}{\sqrt{I(U;U^{\prime})I(V;V^{\prime})}},$ (3)
and VI becomes
@@
The generalization leaves NMI and VI unchanged for hard clustering, where $I(U;U^{\prime})=H(U)$, and enables the measures to be applied to both soft and hard assignments.
@@
Finally, we note that the generalized VI is no longer a proper metric, as the triangle inequality is not guaranteed (Appx. C).
Figure 2: Comparing similarity measures for synthetic embedding spaces. Left: A dataset of 64 points, $x=1,...,64$, is transformed into nine representation spaces marked a-i. Each posterior distribution $p(u|x)$ is a Gaussian with diagonal covariance matrix and standard deviations indicated by the colored ellipses. Right: Pairwise similarity (blue matrices) and pairwise distances (red matrices) for generalized NMI, VI, and CKA, derived using either Monte Carlo or the Bhattacharyya fingerprint, and stochastic shape metrics from Duong et al. (2023). Runtimes to calculate the full matrix are shown above each method. The stochastic shape metric requires the dimensionality of the compared spaces to match; undefined entries comparing a 2D space to space f are marked with an X.
3.2 Quick estimation based on fingerprinting a representation space
Mutual information can be challenging and costly to estimate (McAllester & Stratos, 2020), and we will often be interested in computing all pairwise similarities between many spaces. Fortunately our primary interest lies in scenarios that simplify the estimation considerably: we will focus exclusively on representation spaces whose members are Gaussian distributions, and often the points of comparison will be individual channels (dimensions) of a VAE or GAN latent space that convey only a few bits of information.

### v3

@@ Abstract
Probabilistic representation spaces convey information about a dataset and are shaped by factors such as the training data, network architecture, and loss function. Comparing the information content of such spaces is crucial for understanding the learning process, yet most existing methods assume point-based representations, neglecting the distributional nature of probabilistic spaces. To address this gap, we propose two information-theoretic measures to compare general probabilistic representation spaces by extending classic methods to compare the information content of hard clustering assignments. Additionally, we introduce a lightweight method of estimation that is based on fingerprinting a representation space with a sample of the dataset, designed for scenarios where the communicated information is limited to a few bits. We demonstrate the utility of these measures in three case studies. First, in the context of unsupervised disentanglement, we identify recurring information fragments within individual latent dimensions of VAE and InfoGAN ensembles. Second, we compare the full latent spaces of models and reveal consistent information content across datasets and methods, despite variability during training. Finally, we leverage the differentiability of our measures to perform model fusion, synthesizing the information content of weak learners into a single, coherent representation. Across these applications, the direct comparison of information content offers a natural basis for characterizing the processing of information.
@@ 1 Introduction
The comparison of representation spaces is a problem that has received much attention, particularly as a route to a deeper understanding of information processing systems (Klabunde et al., 2023; Mao et al., 2024; Huh et al., 2024; Lin & Kriegeskorte, 2024) and with applications in tasks like transfer learning, ensembling, and other forms of representational alignment (Sucholutsky et al., 2023; Muttenthaler et al., 2024). Existing methods are applicable to point-based representation spaces, including centered kernel alignment (CKA) (Kornblith et al., 2019) and representational similarity analysis (RSA) (Kriegeskorte et al., 2008). For representation spaces whose citizens are probability distributions, such as in variational autoencoders (VAEs) or in biological systems with inherent stochasticity, failure to account for the distributed nature of representations can miss important aspects of the relational structure between data points (Duong et al., 2023). Few methods account for the distributional nature of representations (Klabunde et al., 2023).
One of the few existing methods for comparing probabilistic representation spaces, stochastic shape metrics (Duong et al., 2023), relies on geometric assumptions and fixed transformations, which can limit its flexibility. In contrast, we adopt an information-theoretic approach that evaluates representation spaces based on the information they transmit, providing a method that is agnostic to properties like dimensionality or whether the spaces are discrete or continuous. Specifically, we compare two probabilistic representation spaces using quantities derived from the mutual information between them, capturing their relational structure while ensuring broad applicability across diverse contexts.
We take as a motivating example the task of unsupervised disentanglement, whose goal is to break information about a dataset into useful factors of variation (Higgins et al., 2018; Locatello et al., 2019; Träuble et al., 2021; Balabin et al., 2023; Van Steenkiste et al., 2019). As an example, a representation space might be trained on images of cars so that color, orientation, and model information are separated into different latent dimensions without any supervision about such factors. When ground truth factors of variation are unavailable for evaluation, as is generally the case for real-world datasets, existing evaluation methods assess the degree of consensus in an ensemble of trained models (Duan et al., 2020; Lin et al., 2020). However, the relatedness of representation spaces has failed to account for the probabilistic nature of the representations, reducing posterior distributions to their means and then using point-based comparisons such as correlation (Duan et al., 2020). With a direct comparison of the information content of representation spaces, we stand to improve the characterization of consensus and of unsupervised disentanglement more generally.
@@
In this work, we generalize measures of similarity of the information content of hard clustering assignments (Fig. 1a) to be applicable to probabilistic representation spaces (Fig. 1b). To motivate the generalization, we treat probabilistic representation spaces as soft clustering assignments of data, whereby partial distinguishability between data points is expressed by the overlap between posterior distributions. We then assess the information content of learned representation spaces produced by models in an ensemble of random initializations, and study the effects of method, training progress, and dataset on information processing.
@@
Figure 1: Similarity of representation spaces. In this work, we generalize measures to compare the information content of clustering assignments to apply to probabilistic representation spaces. (a) A hard clustering assignment, such as the living/non-living distinction conveyed by clustering $V$, communicates certain information about the dataset (here, CIFAR-10 images). Comparing the information content of different clustering assignments enables comparative analyses between algorithms, model fusion, and benchmarking. (b) We generalize measures for comparing hard clustering assignments to be applicable to probabilistic representation spaces, by recognizing the latter as soft clustering assignments. When cast in terms of information content, there is no requirement for the dimensionality of the spaces to match, and hard clusterings (e.g. labels or annotations) can be compared to probabilistic spaces.
@@ 2 Related work
The proposed method builds upon classic means of comparing the information content of different cluster assignments (clusterings) of a dataset and generalizes them to compare probabilistic representation spaces. We then use the method to empirically study generative models designed to fragment information about a dataset in a learned latent space, i.e., for unsupervised disentanglement.
@@
Similarity of clusterings and of representation spaces. The capacity to compare transformations of data produced by different machine learning models enables ensemble learning, a deeper understanding of methodology, and benchmarking (Punera & Ghosh, 2007). Strehl & Ghosh (2002) used outputs of clustering algorithms to perform ensemble learning, based on a measure of similarity between clustering assignments that we will extend in this work: the normalized mutual information (NMI). Referred to as consensus clustering or ensemble clustering, efforts to combine multiple clustering outputs can leverage any of a variety of similarity measures (Wagner & Wagner, 2007; Vinh et al., 2009; 2010; Vega-Pons & Ruiz-Shulcloper, 2011; Huang et al., 2017). Another measure backed by information theory is the variation of information (VI), coined by Meilă (2003) for clustering but recognized as a metric distance between information sources at least twice before (Shannon, 1953; Crutchfield, 1990).
@@
Commonly referred to simply as ‘clustering’, hard clusterings assign every input datum to one and only one output cluster and have been generalized to multiple forms of soft clustering (Campagner et al., 2023). We focus specifically on fuzzy clustering, where membership is assigned to multiple clusters and must sum to one for each datum (Zadeh, 1965; Dunn, 1973; Ruspini et al., 2019). We observe that a probabilistic representation space communicates for each datum a soft assignment over latent vectors, with the degree of membership expressed by the probability density of posterior distributions. Comparisons of hard clustering have been extended to fuzzy clustering for measures that indirectly assess information content, such as the Rand index (Punera & Ghosh, 2007; Hullermeier et al., 2011; D’Ambrosio et al., 2021; Andrews et al., 2022; Wang et al., 2022). Information-based measures have been extended for specific types of soft clusters (Campagner & Ciucci, 2019; Campagner et al., 2023) but none, to our awareness, for comparing the information content of fuzzy clusterings over a continuous representation space.
@@
A rich area of research compares point-based representation spaces via the pairwise geometric similarity of a common set of data points in the space (Kornblith et al., 2019; Hermann & Lampinen, 2020; Klabunde et al., 2023), building upon representational similarity analysis from neuroscience (Kriegeskorte et al., 2008). In place of geometric similarity, topological similarity more closely probes the information available for downstream processing by employing specific similarity functions with tunable parameters (Lin & Kriegeskorte, 2024; Williams, 2024). For comparing probabilistic representation spaces, stochastic shape metrics (Duong et al., 2023) extend a distance metric between point-based representation spaces based on aligning one with another through prescribed transformations (e.g., rotations) (Williams et al., 2021). By contrast, the information-theoretic lens we adopt requires no enumeration of transformations nor tuning of parameters, and directly assesses the information available for processing by downstream neural networks.
@@
Unsupervised disentanglement. Disentanglement is the problem of splitting information into useful pieces, possibly for interpretability, compositionality, or stronger representations for downstream tasks. Shown to be impossible in the fully unsupervised case (Locatello et al., 2019; Khemakhem et al., 2020), research has moved to investigate how to utilize weak supervision (Khemakhem et al., 2020; Sanchez et al., 2020; Vowels et al., 2020; Murphy et al., 2022) and incorporate inductive biases (Balabin et al., 2023; Chen et al., 2018; Rolinek et al., 2019; Zietlow et al., 2021; Hsu et al., 2024).
@@
A significant challenge for disentanglement is evaluation when no ground truth factorization is available. While one proposed route to evaluation relies on a characterization of the posterior distributions in a model (PIPE, Estermann & Wattenhofer (2023)), a more common approach assesses consensus among trained models in an ensemble of randomly initialized repeats. The motivation, first proposed in Duan et al. (2020), is that disentangled models are more similar to each other than are entangled ones because there are intuitively more ways to entangle information than to disentangle it. There, the similarity between two models was evaluated with an ad hoc function of dimension-wise similarity, which was computed as the rank correlation or the weights of a linear model between embeddings in different one-dimensional spaces. ModelCentrality (Lin et al., 2020) quantified similarity between models with the FactorVAE score (Kim & Mnih, 2018) where one model’s embeddings serve as the labels for another model. However, none of these methods fully account for the distributional nature of representations, instead relying on point estimates such as posterior means or single embeddings. In contrast, our approach focuses on capturing the probabilistic structure of representation spaces, and additionally shifts the emphasis from model similarity to channel (or latent subspace) similarity under the premise that the fragmentation of information, central to disentanglement, is more naturally studied via the information fragments themselves.
@@ 3 Method
Our goal is to compare the information transmitted about a dataset by different representation spaces, and we will use the comparison of hard clustering assignments (e.g., the output of $k$-means) as a point of reference. Analogously to hard clustering (Fig. 1a), probabilistic representation spaces communicate a soft assignment over embeddings that is expressed by a probability distribution in the space for each data point (Fig. 1b).
While the proposed method can be applied to any representation space with probabilistic embeddings, we will focus primarily on variational autoencoders (VAEs) (Kingma & Welling, 2014). Let the random variable $X$ represent a sample $x\sim p(x)$ from the dataset under study. $X$ is transformed by a stochastic encoder parameterized by a neural network to a variable $U=f(X,\epsilon)$, where $\epsilon$ is a source of stochasticity. We will make use of two fundamental quantities in information theory (Cover & Thomas, 1999), the entropy of a random variable $H(Z)=\mathbb{E}_{z\sim p(z)}[-\log p(z)]$ and the mutual information between two random variables $I(Y;Z)=H(Y)+H(Z)-H(Y,Z)$. The encoder maps each data point $x$ to a posterior distribution in the latent space, $p(u|x)$, after which a point $u\sim p(u|x)$ is sampled for downstream processing. Commonly, the posterior distributions are parameterized as normal distributions with diagonal covariance matrices, which will facilitate many of the involved measurements but is not required for the method.
@@ 3.1 Comparing representation spaces as soft clusterings
Consider a hard clustering of data as communicating certain information about the data (Fig. 1a). By observing the cluster assignment $U$ instead of a sample $X$ from the dataset, information $I(X;U)$ will have been conveyed. For a hard clustering, every data point is assigned unambiguously to a cluster—i.e., $H(U|X)=0$—which makes the communicated information equal to the entropy of the clustering, $I(X;U)=H(U)$. We note that maximizing communicated information, such as by assigning each data point to its own cluster, does not yield a useful representation. The value of a representation lies in the balance between the information preserved and the irrelevant variation discarded, highlighting the importance of assessing its specific information content.
@@
Soft clustering extends hard clustering to allow each datum to have partial membership in multiple clusters, allowing partial distinguishability between data points to be communicated (Zadeh, 1965; Ruspini et al., 2019) (Fig. 1b). While soft clustering is predominantly performed over a discrete set of clusters, here we view each point $u$ in a continuous latent space as a cluster, with the posterior distribution $p(u|x)$ assigning membership over the continuum.
The utility of NMI or VI over raw mutual information largely resides in the standardization of values. For two identical hard clustering assignments, $U$ and $U^{\prime}$, $\text{NMI}(U,U^{\prime})=1$ and $\text{VI}(U,U^{\prime})=0$. In contrast, soft assignments include additional entropy stemming from uncertainty that is unrelated to the information communicated about the dataset, as is evident from the relation $H(U)=I(X;U)+H(U|X)$ with $H(U|X)>0$. As a result, the entropy terms in Eqns. 1 and 2 that ground the mutual information $I(U;V)$ have components that disrupt the standardization of values. To address this issue, we propose replacing the entropy of a clustering assignment with the mutual information between two copies of that assignment: specifically, substituting $H(U)$ with $I(U;U^{\prime})$, and similarly for $V$. For hard clustering, $H(U)=I(X;U)=I(U;U^{\prime})$, making these quantities interchangeable. Only the third quantity, $I(U;U^{\prime})$, maintains the standardization of $\text{NMI}=1$ and $\text{VI}=0$ for identical soft clustering assignments. The generalized forms then become
$\text{NMI}(U,V)=\frac{I(U;V)}{\sqrt{I(U;U^{\prime})I(V;V^{\prime})}},\ \text{and}$ (3)
@@
The generalization leaves NMI and VI unchanged for hard clustering, where $I(U;U^{\prime})=H(U)$, and enables the measures to be applied to both soft and hard assignments. We note that the generalized VI is no longer a proper metric, as the triangle inequality is not guaranteed (Appx. D).
@@
3.2 Routes to estimation
We propose two routes to estimating NMI and VI that offer a tradeoff between precision and speed; both leverage the known posterior distributions to calculate the information transmitted about the dataset by combinations of representation spaces, $I(X;\cdot)$. The first route is to compute $I(X;\cdot)$ with a straightforward Monte Carlo estimate using the aggregated posterior over the entire dataset of size $L$,
$I(X;U)=\mathbb{E}_{x\sim p(x)}\mathbb{E}_{u\sim p(u|x)}\left[\log\frac{p(u|x)}{\sum_{i}^{L}p(u|x_{i})}\right].$ (5)